### v1

# Twists, Humps, and Pebbles: The Uneven Landscape of End-to-end Speech Recognition and Demographics

Abstract
Current voice recognition approaches use multi-task, multilingual models for speech tasks like Automatic Speech Recognition (ASR) to make them applicable to many languages without substantial changes. However, broad language coverage can still mask performance gaps within languages, for example, across genders. We systematically evaluate multilingual ASR systems on gendered performance gaps. Using two popular models on three datasets in 19 languages across seven language families, we find clear gender disparities. However, the advantaged group varies between languages. While there are no significant differences across groups in phonetic variables (pitch, speaking rate, etc.), probing the model’s internal states reveals a negative correlation between probe performance and the gendered performance gap. I.e., the easier to distinguish speaker gender in a language, the more the models favor female speakers. Our results show that group disparities remain unsolved despite great progress on multi-tasking and multilinguality. We provide first valuable insights for evaluating gender gaps in multilingual ASR systems. We release all code and artifacts at https://github.com/g8a9/multilingual-asr-gender-gap.
Multilingual Speech Models for Automatic Speech Recognition Exhibit Gender Performance Gaps
Giuseppe ${}^{\clubsuit}$, Beatrice ${}^{\spadesuit}$, Dennis ${}^{\diamondsuit}$${}^{\spadesuit}$, Dirk ${}^{\clubsuit}$$^{\clubsuit}$ Bocconi University, Milan, Italy ${}^{\diamondsuit}$ University of Trento, Trento, Italy ${}^{\spadesuit}$ Fondazione Bruno Kessler, Trento, Italy giuseppe.attanasio3@unibocconi.it
1 Introduction
Recent advancements in Automatic Speech Recognition (ASR) have showcased remarkable capabilities of “one-model-fits-all” multi-task, multilingual solutions (Radford et al., 2022; Communication et al., 2023; Chu et al., 2023). These neural networks can transcribe, translate, or generate speech in an unprecedented number of languages. Consequently, these models have found widespread adoption [^1] Anecdotally, all major cloud providers and countless smaller companies list a tutorial on how to deploy and serve Whisper (Radford et al., 2022) on their platforms. and are being integrated into applications for the general public. [^2] https://techcrunch.com/2023/06/19/gladia-turns-any-audio-into-text-in-near-real-time/
Figure 1: Model error rate gaps between female and male speakers. Average across language families and data sets. Empty shapes show models that advantage women, filled shapes advantage men. Pairs with one shape indicate gap on one group only. Language count per family in brackets. Ideal scores on the “0” line.
However, applicability to many languages should be one of many optimizations. Voice and language production are among the strongest identity traits, widely varying across people, including genders. These differences have physiological and sociocultural roots (e.g., Labov, 1964; Wolfram, 2004; Alim, 2004; Zimman, 2017, 2020). While physiological differences are rooted in biological sex (e.g., vocal tract length), sociocultural differences (e.g., prescribed registers) reflect gender identity. In this paper, we account for both.
Because of their varied features and societal implications, gendered voice and speech differences—usually as binary distinction—have been extensively studied (Coleman, 1976; Busby and Plant, 1995; Hillenbrand and Clark, 2009). However, modeling gender also poses various challenges for ASR, often leading to performance disparities usually favoring male Tatman (2017); Garnerin et al. (2019, 2021) but sometimes female speakers Adda-Decker and Lamel (2005); Feng et al. (2021); Liu et al. (2022). Those performance disparities can lead to unequal service quality, which is especially harmful to already socially disadvantaged groups (Mengesha et al., 2021). However, most of those studies are within one language (usually English). Evaluating the now widely deployed multilingual ASR models frequently overlooks gendered differences, posing fairness concerns.
We are the first to evaluate gendered performance gaps in massively multilingual ASR models systematically. More specifically, we pose two research questions. (Q1) Do multi-task, multilingual ASR models perform equally well across speakers identifying as women, men, or neither of the two? And, if so, (Q2) can we relate performance gasp to (socio)phonetic features or how models encoded gender internally?
To answer Q1, we evaluate two state-of-the-art multilingual ASR models on three datasets (read and spontaneous speech) covering 19 languages from seven language families. Our results show that models do not perform equally well across genders. Whether models favor male or female speakers depends on the dataset and language. Figure 1 shows average error rates across the tested language families. The model performance of speakers who do not identify with either gender is lower than that of those identifying as male. We conduct phonetics and interpretability experiments to understand the reasons for those gaps (Q2). Phonetic analysis shows no significant differences across genders and cannot explain the observed phenomenon. Interpretability through gender probes reveals a negative correlation between female-male gaps and the ability to discern data across genders from a model’s internal representations. This finding suggests that when presented with speech from men and women, models build different internal representations that can serve as a proxy for gender disparities.
Contributions.
We conducted the first extensive evaluation of gender performance differences between two widely used multilingual ASR models and found significant gender gaps. We release code, data, and all artifacts we produce for future research.
Bias Statement.
When using gender as a variable, we focus on the speaker’s declared identity. We evaluate whether speech from individuals identifying as female, male, or neither of the two are equally processed by multilingual ASR models. Conceptually, this is the ideal outcome, and we measure performance disparities of minority groups (“female”, “other”) versus the socially advantaged group, i.e., men. We define a system to be biased if it risks further contributing allocational (i.e., quality of service, technology less accessible) and representational harms impacting the minority groups, e.g., feeding into stereotypes about inadequacy of women and speech technology or “shrill,” “incorrect” voice (Tallon, 2019). [^3] Voice technology expert Tom Schalk once said that “many issues with women’s voices could be fixed if female drivers were willing to sit through lengthy training… Women could be taught to speak louder and direct their voices towards the microphone.” https://techland.time.com/2011/06/01/its-not-you-its-it-voice-recognition-doesnt-recognize-women/
2 Background
Multilingual Speech Models.
Contemporary speech recognition systems based on Transformers (Vaswani et al., 2017) outperform the previous transducer-based systems (Li et al., 2022). Among these new systems, a particularly successful subset uses multi-task, multilingual approaches. These architectures process audio and text with separate Transformers and use different communication mechanisms to enable cross-modal interactions. OpenAI’s Whisper (Radford et al., 2022) uses a standard encoder-decoder architecture. Audio and text sequences can be arbitrary—to differentiate among tasks and languages, it uses special decoding tokens. This strategy allows multi-tasking and multilinguality without architectural changes or fine-tuning (see Figure 1, top-right, in Radford et al. (2022)). Similarly, Meta’s SeamlessM4T (Communication et al., 2023) uses a speech encoder for audio and an encoder-decoder transformer for text (Costa-jussà et al., 2022, NLLB).
While these models have shown strong multilingual capabilities, current assessments primarily focus on overall performance metrics but lack, for example, fairness evaluations in speech setups. [^4] Seamless M4T (Communication et al., 2023) has been previously evaluated for gender bias in translation tasks.
Gender and Speech (Technologies).
Gendered aspects of the voice are one of the most salient individual traits (Kreiman and Sidtis, 2011; Azul, 2015; Zimman, 2021) and have consequently long been studied in linguistics (Zimman, 2020). The anatomy and makeup of the vocal tract do play a role, as it determines pitch range and formant variations, often regarded as the most distinctive phonetic vocal features of (cis)gender men and women. [^5] Cisgender refers to individuals whose gender identity corresponds to the sex assigned at birth (Fuchs and Toda, 2010). However, gender variation in the voice has been shown to arise from several aspects besides physical ones (Oates and Dacakis, 2015; Zimman, 2018; Becker et al., 2022). As studies including transgender individuals show Zimman (2017), sociocultural factors influence vocal use, including which parts of the pitch range are used (Loveday, 1981; Yuasa, 2008), and articulatory practices for sibilant consonants and vowels that are perceived as gender characterizing (Pharao et al., 2014; Podesva et al., 2016; Li, 2017).
For speech technologies, sociodemographic variation has posed challenges to ASR systems (Sawalha and Abu Shariah, 2013; Liu et al., 2022; Rajan et al., 2022; Fucci et al., 2023). Various works have found ASR models for English and French to better recognize male speech and voices (Tatman, 2017; Garnerin et al., 2019, 2021). This effect is often a result of the under-representation of women in the training data (Meyer et al., 2020; Gaido et al., 2020; Garnerin et al., 2020). However, various works have found the reverse to be true. Several studies found comparatively better performance for women in more spontaneous, conversational data in Dutch (Feng et al., 2022) and English (Adda-Decker and Lamel, 2005; Koenecke et al., 2020). They attribute these findings to sociolinguistic factors, like the higher incidence of disfluencies and informal speech in men.
Thus, the current literature provides a fragmented picture of gender disparities in speech technologies. However, most studies focus primarily on a single language, usually English, albeit across multiple datasets and models. We broaden this research to include large-scale multilingual speech models. In addition, we evaluate performance in a third gender category, which includes individuals who do not identify as male or female.
3 Methodology
This section describes the experimental design (§3) to answer Q1, i.e., whether there are gendered performance differences in multi-task multilingual ASR models. Our results (§4) confirm it.
Models.
We experiment with OpenAI’s Whisper (Radford et al., 2022) and Meta’s SeamlessM4T (Communication et al., 2023), two widely used, state-of-the-art multilingual ASR models. See Appendix A.1 for implementation details.
Datasets.
Among other multilingual datasets available for ASR (Gales et al., 2014; Black, 2019; Pratap et al., 2020; Iranzo-Sánchez et al., 2022; Valk and Alumäe, 2021), we use Mozilla Common Voice (Ardila et al., 2020, CV), [^6] We use CV 16.0 from https://huggingface.co/datasets/mozilla-foundation/common_voice_16_0 Fleurs (Conneau et al., 2023) and VoxPopuli (Wang et al., 2021, VoxPopuli). Our choice of datasets is bound to the availability of reliable speakers’ gender information. [^7] Gender is self-declared in CV and Fleurs, and retrieved from https://multimedia.europarl.europa.eu/en in VoxPopuli. Furthermore, these datasets cover various recording and speech conditions: i) read, where speech is typically well-articulated and based on the reading of pre-defined texts (CV and Fleurs), and ii) spontaneous conditions elicited from public talks and speeches, which allow for more speaker-dependent articulations and variation in word usage (Gabler et al., 2023, VoxPopuli). The type of recording conditions might influence the size of performance disparities, with spontaneous speech providing a more realistic image of the impact of potential disparities for real-use ASR cases.
To avoid unreliable results due to data contamination, we base our analysis on the concatenation of validation and test splits of each dataset. See Appendix A.4 for a discussion on data contamination and how transcription performance varies across splits in our setup.
Languages.
We include 19 languages from CV and Fleurs, and a subset of those—11 in total, due to data availability—from VoxPopuli. These languages represent diverse language families and data availability conditions (i.e., high-low resource). Our language selection depended on sufficient utterances stratified across genders to allow for proper comparisons. See Table 4-6 for full statistics on datasets in Appendix A.5.
3.1 Evaluation
Quality Metrics.
We use standard Word- and Character-Error Rate (WER and CER, respectively) to evaluate transcription quality. Following Radford et al. (2022), we i) report WER for all languages but Yoruba and Japanese—where variability in orthography and ambiguous word units may affect evaluation Rowlands (1954); Matsuoka et al. (1997)—for which we use CER, ii) normalize all references and hypotheses texts by transliterating them into Cyrillic for Russian and Serbian, and iii) when evaluating Whisper, we apply the official normalization routines to texts. [^8] Found at https://github.com/openai/whisper/tree/main/whisper/normalizers Moreover, we trim audio to the initial 30 seconds  [^9] Most of the recordings in our datasets are shorter than 30 seconds. See statistics in Table 4-6 in Appendix A.5. and filter out dataset noise by removing records with an empty reference and silence in the snippet. [^10] When processing Fleurs-es, we noticed some records containing only background noise. We used state-of-the-art voice activity detection models to detect and filter them out. See Appendix A.2 for details.
When evaluated in terms of overall quality, both Whisper and Seamless showcase competitive results, respectively, on: Fleurs (12.68 and 19.87 – avg. 19 langs); CV (17.51 and 16.42 – avg. 19 langs); VP (13.59 and 12.74 – avg. 11 langs). For disaggregated WER scores per each language and dataset, see Figure 7 in Appendix A.1.
Gap Metrics.
Measuring the ASR gender gap is equivalent to seeking what is commonly known in the field of machine learning as group fairness (Chouldechova and Roth, 2020), or, more specifically, demographic parity (Dwork et al., 2012). Conceptually, parity is achieved when a given statistical measure is equal across different groups.
We operationalize demographic parity using the notation from Czarnowska et al. (2021). We introduce a Pairwise Comparison Metric, loosely inspired by the Disparity Score in Gaut et al. (2020), defined as follows:
$E(r_{A},r_{B})=100\cdot(\phi(r_{A})-\phi(r_{B}))/\phi(r_{B})$ (1)
where $r_{A}$ and $r_{B}$ are the set of audio snippets belonging to given groups (e.g., either “male”, “female”, or “other” in CV), and $\phi(\cdot)$ is one of the quality metrics defined above (i.e., WER or CER). The ideal score is 0—i.e., the model performs perfectly equally on the two groups. [^11] ASR models achieve different baseline error rates across languages. Equation 1 measures relative, rather than absolute, gaps to account for this variability. See Table 3 in Appendix A.4 for details on absolute gaps and §7 for a discussion on possible metric interpretations.
Significance Test.
We use a bootstrapped approach (Søgaard et al., 2014) to estimate $\phi(\cdot)$. We run $n=1000$ iterations sampling 40% of the set. When sampling, we balance the two groups in several records for a reliable comparison. We compute $E(\cdot)$ on the arithmetic means of $\phi(\cdot)$ across the $n$ runs and use a two-sided Student’s t-test to compute the statistical difference between the means.
Filtering and bootstrap sampling improve the evaluation robustness of the skewed speaker and gender distributions. See Appendix A.4 for more details on comparing sampling performance vs. complete sets.
Note that sampling requires attributing each record to an individual speaker—information not available in Fleurs. For consistency, we attribute a speaker ID to each record automatically. Specifically, we i) use state-of-the-art speaker verification embedding models to encode each recording, ii) cluster them using HDBSCAN (Campello et al., 2013), and iii) attribute a speaker ID to each cluster. See Appendix A.3 for details on the pipeline.
4 Gender Performance Gaps
In the following, we report the results of evaluating Whisper and Seamless, focusing on gender performance disparities. We first investigate whether models are equally able to recognize speech for the female and male group (§4.1), then for speakers that identify as neither of the two comparatively to the male gender group (§4.2).
(a) VoxPopuli
(b) Fleurs
(c) CV
[…]
4.1 “Female” – “Male” Gap
Figure 2 reports $E(r_{F},r_{M})$, i.e., the gender gap results for Whisper and Seamless on each dataset. Concerning our Q1, a bird’s eye view immediately attests that multilingual models do not perform equally across female and male speakers, showing a preference for one gender over the other for most instances. Our analysis does not consistently reveal models disadvantaging the feminine group—see values below 0. This finding stands in stark contrast with mounting evidence of a strong masculine bias affecting a wide range of Natural Language Processing (NLP) tasks (e.g., Sun et al., 2019; Stanczak et al., 2022). It is especially noteworthy given the well-known under-representation of women in current resources used for model training (Garnerin et al., 2019; Zanon Boito et al., 2022; Sun and Peng, 2021). We observe that both speech models—albeit to varying degrees—better recognize female speech on average in two out of the three considered datasets (i.e., Fleurs and VP).
Upon closer examination, we also observe quite different behaviors across datasets. First, despite the higher degree of spoken language variation to be expected for the spontaneous VP dataset (Figure 1(a)), the evaluation of this elicited resource shows a comparatively reduced gender gap, with overlapping values between Whisper and Seamless for most languages (i.e., hu, es, fr, nl, cs, de). Conversely, it is on read resource that we attest to a higher degree of variability across models as well as languages, with occasional wide performance disparities toward either females or males in CV (1(c)). Notably, Seamless remains closer to performance equality for most instances on Fleurs (1(b)), i.e., values between 11 and -10.
As such, in line with previous research (see §2), analysis of gender disparities for the ASR reveals a complex picture, which we further analyze and unpack in the following sections.
4.2 “Other” – “Male” Gap
ca de en es fr
Whisper 4.68 37.20 28.50 -2.00 5.21
Seamless 52.76 17.19 38.70 11.37 -0.50
Table 1: Model error rate gap (Eq. 1). Positive values indicate better performance on men, negative ones on “other.” $p<0.05$
Restricting our analysis to validation and test sets leaves us with few records for speakers who do not identify as neither male nor female. As such, we report results for five high-resource languages, Catalan, German, English, Spanish, and French. [^12] Size varies from 16 (ca min.) to 86 (es, max) records
Table 1 reports $E(r_{O},r_{M})$ for both models. Results suggest that models penalize non-male speech and that gaps are larger than disparities observed for the female/male groups. Given the limited sample size, we deem this analysis only exploratory. However, it is a valuable first overview of the model’s behaviors for gender-non-conforming voices. We underscore the need for more representation of diverse voices and sociodemographic groups in current resources, and for which monitoring fairness in existing spoken language models remains out of reach. See §7 for an expanded discussion.
5 Acoustic Analysis
Figure 3: Acoustic features. Speaking rate, pitch average and intensity average (left to right). Italian (2(a)) and Slovak (2(b)) on our three datasets. “Male” (green), “Female” (red), and “Other” (blue, available for CV).
Our findings (§4) reveal performance gaps across gender groups, with variable trends over languages and datasets. To further investigate potential reasons, we explore whether phonetic features, as measured across different groups in our evaluation records, shed light on the phenomenon (Q2). For each segment, we examine three parameters: i) pitch average, the mean of the fundamental frequency values measured in Hertz (Hz) Hirst and de Looze (2021); ii) intensity average, the mean of the fundamental frequency values measured in decibel (dB) for each sample Pausewang Gelfer and Young (1997); iii) speaking rate, computed as the number of tokens per minute Künzel (2013). [^13] We compute average pitch and intensity using Praat Boersma and Weenink (2001). Concerning speaking rate, we found no reliable multilingual tools for syllable segmentation. Therefore, we approximated it by the number of tokens as computed by Whisper’s tokenizer.
We analyze these features because they represent sociophonetic variability of interest for the gender dimension and within our setting.
As already discussed in §2, differences in pitch between men and women are well documented Coleman (1976); Hillenbrand and Clark (2009), and variations in speaking rate have also been associated with gendered speech Borsel and Maesschalck (2008). Besides, the selected features might capture distinctive acoustic aspects across the datasets used in our experiments. For instance, intensity can be influenced by the recording conditions Maryn and Zarowski (2015), which can vary significantly in crowdsourced speech data. Additionally, speech rate tends to be higher in spontaneous speech Nakamura et al. (2008). Differences in acoustic parameters may also exist across languages. For example, languages can vary in the average speaking rate Coupé et al. (2019) and pitch Loveday (1981); Yuasa (2008).
Findings.
Since we found similar distributions across languages – and due to space constrains – in Figure 3 we report results for two representative languages only: Italian and Slovak. The former exemplifies the set of languages with comparable performance across genders, whereas the latter is biased towards men and belongs to the set of languages with larger gender gaps trends (§4).
By comparing these two groups of languages, pitch emerges at the only feature that differ across female and male, whereas no substantial variation is attested for intensity and speaking rate. Interestingly, the “other” group is slightly more spread on pitch and intensity compared to “male” and “female”. By comparing datasets, we still find similar patterns, except for the intensity parameter. Intensity values are more divergent in their distributions, which is particularly noticeable for Fleurs. [^14] This could be attributed to the underlying recording conditions of Fluers, details of which are however not provided.
Overall, however, we do not attest to significant acoustic differences across datasets and languages, regardless of the gap size or direction (male, female, other). Thus, these findings suggest that acoustic features cannot explain gender-based performance gaps and their trends.
6 Probing Gender in ASR Models
Since models exhibit gender-based disparities and acoustic features do not fully explain the phenomenon, we turn to interpretability analysis to understand whether models encode records differently across genders.
The field of Natural Language Processing (NLP) has now established that transformer language models encode syntactic (Hewitt and Manning, 2019), semantic (Tenney et al., 2019), and factual (Petroni et al., 2019; Meng et al., 2022) information in hidden representations. As such, recent work has focused on extracting this information through probes, i.e., supervised classifiers trained on the model’s embeddings (Alain and Bengio, 2016; Belinkov and Glass, 2019). Compellingly, some have related extractability of sensitive attributes, e.g., gender, to bias in downstream application (Orgad et al., 2022). Probing is primarily motivated by the risk that models entangle protected attributes and predictions in sensitive use cases (Zhao et al., 2018; Ravfogel et al., 2020).
Motivated by this line of research, we measure gender extractability in ASR models and ask (Q2) whether and to which extent it explains gender-based performance gaps. Ours is the first study of this kind. [^15] Interpretability for speech models is a relatively new research avenue. Mohebbi et al. (2023) use context-mixing techniques to explain homophone disambiguation; Pastor et al. (2023) use word-segment masking to explain semantic classication tasks.
Experimental design.
We focus on one dataset-model configuration, namely Whisper and CV. We attach our probes to the model encoder’s last layer embeddings and train one distinct probe for every position. We use Logistic Regression and Minimum Description Length (MDL) (Voita and Titov, 2020) to probe gender in the female-male binary setup (see Appendix A.6 for details).
Figure 4: F1 Macro (top) and measured code length ($1e^{6}$, bottom) for Logistic Regression and MDL probes, respectively. English, F-M setup, first 10 seconds of context. Actual score (pale line), exponential moving average (solid), and average length of test snippets (vertical dashed).
Findings.
Figure 4 shows gender extractability on the English CV split for the F-M setup. Trends are similar across languages (full results in Figure 8(a)). First, F1 scores suggest that gender extractability is relatively easy. Whisper produces representations of female and male speech that can be easily separated. Second, extractability is not constant over time. It starts high in the first milliseconds, drops during the actual signal (i.e., the speaker is talking), and finally plateaus around the initial value. This finding recalls attention “sink” theory Xiao et al. (2023) implicating that transformers prioritize initial positions for specific tasks. Third, MDL probes share similar trends—low codelength indicates an “easier” task / higher extractability—suggesting that simple logistic regression is a robust method to start with, countering issues found instead in NLP (Hewitt and Liang, 2019; Voita and Titov, 2020). [^16] To confirm this hypothesis, we conduct additional experiments by training logistic probes with randomly shifted labels. We report random guess performance on all languages (see Figure 8(b)) in Appendix A.6.
Correlation with Error Rate.
[…]
7 Discussion
Our extensive analysis confirms that multi-task multilingual ASR models display gender performance gaps. Depending on the language, disparities can favor male or female speakers, but rarely those who do not identify as either. In the following, we report two discussion points to help frame our study, its findings, and its implications.
Figure 6: Whisper distribution of error rates. Female and male records in Fleurs. nl (top) and ar (bottom). Proportion of records (y axis) having a WER error rate (x axis) at least that large.
Interpreting $E(\cdot)$.
For our gender gap analysis, we study performance disparities by studying the relative difference ($E(\cdot)$) in scores across gender groups (see §3.1). While such definition allows to compare models and languages at different absolute error levels, relative differences might tell little about the concrete impact of gender disparities in ASR. [^18] For instance, a 10% F-M relative difference corresponds to 0.2 or 2 WER points gap if we consider a model that scores 2 or 20 absolute WER on men, respectively. Ideally, a more informative index would indicate the usability of an ASR system, thus providing a more precise estimation of its failures and their severity.
Sentence-level error metrics can provide such a fine-grained view. Following Koenecke et al. (2020), in Figure 6, we report the proportion of generated transcriptions whose quality is greater than given WER values. Due to space constraints, we only show the results for Whisper on Fleurs for two languages as exemplars (i.e., Arabic and Dutch). This visualization lets us identify where performance between genders starts to diverge. Let us hypothesize that transcriptions became unusable after WER>20. The plot can help us spot that men and women would be equally affected in Dutch. However, 50% of Arabic transcripts from men would be unsuitable compared to 40% of women’s. Parity would be achieved if we lose the constraint up until around WER=40.
We echo the literature that urges the need for more informative/concrete bias measurement (Delobelle et al., 2022; Orgad and Belinkov, 2022) and advocate measures that can be used to guide model design and inform its reliable deployment. Thus, this evaluation is particularly relevant and a more realistic estimate of the impact of sociodemographic disparities in real-world settings.
Group-Aware Dataset Splits.
Numerical and statistical studies like ours heavily depend on quality and quantity of evaluation data. In this specific case, we would want to have representative splits, i.e., records from many speakers from each group. However, current dataset releases are drastically skewed. For instance, in most of our datasets and language configurations, men are overrepresented in the number of records and speakers. While downsampling can partially solve the former at evaluation time, little can be done with the latter. Moreover, we point out that stratification over gender and speakers to create splits (train/validation/test) affects languages differently. In a high-resource language, an 80/10/10 division still contributes many speakers from all groups in validation and testing. However, in a low-resource setup, only a few speakers are included—in CV most languages counted only one to 20 speakers considering men and women, and frequently none or less than five speakers for “other”.
We argue that dataset releases should factor in the difficulty in collecting samples from less represented groups and provide group-aware splits to facilitate fairness research.
8 Conclusion
We conducted the first extensive evaluation on gender-based performance gaps of Whisper (Radford et al., 2022) and SeamlessM4T (Communication et al., 2023) for ASR. These models consistently exhibit gender bias across 19 languages from seven language families. Depending on the language, disparities can favor men or women but rarely favor speakers who identify as neither. We locate a potential source of these gaps using gender probes from interpretability approaches. Our results show that probes can be a proxy for gender gaps and that group fairness in multi-task multilingual models remains unsolved.
Ethical Considerations
The use of gender as a variable in this warrants some ethical reflection.
As one of the most salient perceptual traits of one’s identity (Kreiman and Sidtis, 2011; Azul, 2015; Zimman, 2021) gendered differences represent a linchpin of much (socio)phonetic research (Zimman, 2020), which has unpacked several physical and sociocultural factors contributing to such gendered characteristics in the voice. Based on this evidence, our work does not intend to be normative nor assumes the existence of a single, unidimensional “female” or “male” voice. Instead, we question whether different gendered groups are equally recognized by current multilingual ASR models and incorporate sociophonetic knowledge in our analysis and discussions to isolate why that might not be the case. To do so, we do not make any inference about the gender of the speakers in the employed data. Instead, we rely on the declared gender of the speakers in the employed speech resources. In this regard, part of our analysis of the Mozilla Common Voice dataset uses a third category, “Other”, which potentially aggregates diverse identities (e.g., transgender, non-binary, and other marginalized individuals) [^19] These are among the offered gender options that can currently be reported when donating one’s voice for Common Voice. under one single umbrella term. While this third category also allows us to include genders non-conforming to the binary in our study, we recognize that this label and category might be an oversimplification and risk erasing the experiences and representativity of many gender identities.
Finally, gender probes represent a methodological approach to studying how models encode different audios from input signals. It is, hence, essential to remember the inherent limitations and ethical implications of relying on probes for gender classification. While these tools may offer a convenient means of categorization and can be suitable for exploring the models’ behavior, they often overlook the nuanced and multifaceted nature of gender identity.
Limitations
Our paper comes with a series of limitations. We divide them into two categories: data and methods.
Data.
Gender is the driving variable of our analysis. However, it is widely recognized that gender interacts with other sociocultural factors, e.g., dialect (Wolfram, 2004) or sexual orientation (Zimman, 2013), in voice production. By limiting to self-identified gender provided with the datasets, our analysis can only provide a partial view. We thus advocate for speech dataset releases that include a principled set of such factors.
The scarcity of data limits the generalizability of our results in two more setups in our analysis. First, the category “Other” counts a deficient number of speakers and records in most of our setups (see §7). Second, the number of speakers we extracted from Fleurs is compared to CV and VoxPopuli (see 3.1). These factors hamper cross-dataset comparison and overall generalizability.
We did not control for data quality on references. It might be that references are noisy (e.g., single words, empty) and can lead to over- or under-estimation of our measurements.
Finally, data contamination. Although we chose validation and test sets as our analysis targets, we cannot exclude that models were trained on part or all of them. We conducted a side analysis (Appendix A.4) that suggested that this might not be the case.
Methods.
Since we miss reliable multilingual phonetic tools, we estimated the speaking rate using Whisper’s pre-trained tokenizer. We acknowledge that pretrained tokenizers yield fewer tokens for high-resource languages. Hence, our measurement of low-resource might be overestimating the phenomenon.
If framed in the context of bias evaluation paradigms—which we are currently not doing—probing can be seen as an intrinsic bias paradigm. We correlate probing performance to a downstream task (and potential harm), i.e., quality gaps in ASR. However, studies in the field of NLP recognize that intrinsic and extrinsic (downstream) bias metrics do not necessarily correlate (Goldfarb-Tarrant et al., 2021; Kaneko et al., 2022). To our knowledge, ours is the first study to inspect such aspects in transformers for speech, and findings may not hold across modalities. We leave this research question to future work.
Appendix A Experimental Details
(a) VoxPopuli
(b) Fleurs
(c) Common Voice
Figure 7: Seamless and Whisper transcription quality. WER results are computed on test splits.
A.1 Multilingual ASR Models
For both Whisper and SeamlessM4T, we used code and model checkpoints in transformers (Wolf et al., 2020). The Hub’s model IDs are openai/whisper-large-v3 and facebook/seamless-m4t-v2-large, respectively. Both checkpoints correspond to the latest and best-performing versions available at the time of writing, February 2024.
We used each models’s standard decoding configurations to transcribe each audio snippet.
In figure 7, we report WER results for Seamless and Whisper on the languages and datasets used in our experiments.
test train val gender set gender subset
Fleurs Whisper 12.68±.${\pm.11.3}$ 12.64±.${\pm.11.4}$ 12.49±.${\pm.11}$ 12.60±.${\pm.11.2}$ 12.62±.${\pm.11.2}$
S 4MT 19.87±.${\pm.6.4}$ 19.72±.${\pm.6.3}$ 18.75±.${\pm.6.2}$ 19.65±.${\pm.6.3}$ 19.67±.${\pm.6.3}$
CV Whisper ${\pm.17.5}$ 16.65±.${\pm.19.8}$ 17.02±.${\pm.18.7}$ 16.88±.${\pm.19.2}$ 18.45±.${\pm.19.2}$
S M4T 16.42±.${\pm.9.5}$ 12.68±${\pm 9.9}$ 14.7±.${\pm.8.9}$ 13.06±.${\pm.9.6}$ 15.58±.${\pm.9.3}$
VP Whisper 13.59±.${\pm.4.45}$ 12.63±.${\pm.3.79}$ 12.99±.${\pm.4.22}$ 12.69±.${\pm.3.8}$ 13.32±.${\pm.4.2}$
S M4T 12.74±.${\pm.3.27}$ 11.27±.${\pm.2.63}$ 12.33±.${\pm.2.68}$ 11.39±.${\pm.2.6}$ 12.54±.${\pm.2.9}$
Table 2: WER across datasets splits (test/train/validation), all presampled female/male records, and on the sampled gender subset used in our experiments. Results averaged over 19 languages in CV and Fleurs, and 11 for VoxPopuli.
test train val presample_all overall
Fleurs Whisper -1.18±.${\pm.2.9}$0.659±.${\pm.1.8}$0.255±.${\pm.2.6}$0.893±.${\pm.1.7}$0.997±.${\pm.1.9}$
S M4T 0.153±.${\pm.1.9}$0.880±.${\pm.1}$0.257±.${\pm.2.7}$0.735±.${\pm.0.72}$0.809±.${\pm.0.72}$
CV Whisper 1.17±.${\pm.5.7}$0.65±.${\pm.2.7}$ 0.88±.${\pm.5.6}$0.14±.${\pm.3.5}$ 1.06±.${\pm.5.3}$
S M4T 2.01±.${\pm.5.4}$1.22±.${\pm.5.9}$ 0.40±.${\pm.2.7}$ 0.59±.${\pm.2.9}$ 1.05±.${\pm.3.3}$
VP Whisper 0.31±.${\pm.3.5}$0.80±.${\pm.0.97}$0.58±.${\pm.1.9}$0.73±.${\pm.0.9}$0.104±.${\pm.2.2}$
S M4T 0.16±.${\pm.2.8}$0.77±.${\pm.0.54}$0.71±.${\pm.2.1}$0.75±.${\pm.0.6}$0.268±.${\pm.2.2}$
Table 3: Average F-M WER absolute difference across datasets splits (test/train/validation), all presampled female/male records, and on the sampled gender subset used in our experiments. The results are averaged over 19 languages for CV and 11 for VoxPopuli. For Fleurs, we average results over the only 9 languages that comprise both female/male speakers in all splits.
A.2 Voice Activity Detection
While conducting our experiments, we noticed that other than records with an empty reference, datasets can contain also empty audio snippets. More precisely, these recordings do have a signal but it is mostly silence. We found this phenomenon primarily in Fleurs-es, with silence mostly coming from snippets attributed to women.
Therefore, we used pyannote.audio’s pretrained neural models and code for voice activity detection. [^20] https://github.com/pyannote/pyannote-audio After counting the number of segments where voice was detected, we filter out all snippets were no segment was detected. We release the the counts and IDs of silence records in our repository.
A.3 Speaker ID Attribution in Fleurs
Figure 8: UMAP projections of snippets from men (left, 13 speakers found) and women (right, 16 speakers found) in Fleurs-en. Color indicates cluster ID assigned by HDBSCAN.
Speaker IDs are crucial in our pipeline to avoid over- or under-estimating performance gaps due to overly present speakers. Since Fleurs does not provide this piece of information, we attribute it automatically.
Specifically, we encode each snippet using a pretrained ECAPA-TDNN model for speaker verification. We use code from SpeechBrain (Ravanelli et al., 2021) and the model checkpoint with Hub ID speechbrain/spkrec-ecapa-voxceleb. Next, we cluster embeddings separately per gender using HDBSCAN (Campello et al., 2013), with code from scikit-learn. We use set min_cluster_size to 2 and use cosine similarity. Finally, we assign a numerical ID to each found cluster. Figure 8 shows an example of clustered snippets projected with UMAP (McInnes et al., 2018). We release per dataset and language IDs in our repository.
A.4 Subset selection and data splits
To ensure the soundness of our experimental framework and of the gender subsample selection described in in Section 3, we conducted preliminary analyses by comparing WER transcription quality scores (Table 2) and gender F-M gap results (Figure 3) across: i) different datasets splits (test/train/validation), ii) all female/male records available from any split, and iii) the sampled female/male subset used in our main experiments.
Since no precise information concerning the data used to train Whisper and Seamless is publicly available, this evaluation is carried out with the goal of spotting potential data contamination issues. Namely, to ensure that we are not testing models on data samples comprised in their underlying training, thus posing the risk of obtaining unreliable results, which cannot be compared across datasets and models.
Figure 2 shows no notable WER variations for Fleurs. Instead, both models achieve better performance on the train splits of CV and VP, especially for the former dataset. The influence of the split accounted in CV and VP can thus be reflective of the F-M WER differences reported in Figure 3. Informed by such results, the gender subset used in our experiments is sampled from the test and validation split only for CV and VP, whereas for Fleurs we leverage gender data from the whole corpus.
A.5 Dataset Statistics
Lang # records Seconds # Tokens # M Gini (M) # F Gini (F)
ar 14013 4.37±$\pm 1.51$ 19.92±$\pm 17.55$ 319 0.66 102 0.68
ca 1645 6.12±$\pm 1.73$ 17.05±$\pm 6.22$ 265 0.22 241 0.22
cs 13746 4.60±$\pm 1.38$ 20.77±$\pm 8.71$ 293 0.64 39 0.61
de 3641 6.10±$\pm 1.75$ 16.74±$\pm 6.08$ 807 0.22 129 0.24
en 4938 5.96±$\pm 2.28$ 11.08±$\pm 3.99$ 2262 0.25 476 0.25
es 6404 6.09±$\pm 1.60$ 15.61±$\pm 5.01$ 1476 0.29 517 0.30
fi 1695 4.74±$\pm 1.66$ 18.22±$\pm 8.93$ 52 0.57 17 0.57
fr 4214 5.81±$\pm 1.67$ 16.40±$\pm 5.56$ 876 0.19 213 0.19
hu 12245 5.45±$\pm 1.52$ 22.84±$\pm 9.23$ 173 0.57 277 0.53
it 7088 6.14±$\pm 1.69$ 17.52±$\pm 6.44$ 888 0.29 203 0.28
ja 7981 5.02±$\pm 2.29$ 22.64±$\pm 15.07$ 1033 0.44 632 0.41
nl 12808 4.83±$\pm 1.38$ 16.29±$\pm 5.57$ 487 0.62 130 0.63
pt 7319 4.61±$\pm 1.63$ 10.66±$\pm 5.44$ 620 0.41 82 0.37
ro 6489 4.05±$\pm 0.90$ 15.70±$\pm 4.13$ 167 0.56 39 0.57
ru 10466 5.51±$\pm 1.85$ 18.87±$\pm 9.18$ 702 0.45 248 0.43
sk 3008 4.34±$\pm 1.74$ 13.66±$\pm 9.64$ 58 0.64 13 0.65
sr 1895 2.86±$\pm 0.89$ 8.63±$\pm 5.97$ 39 0.57 11 0.64
sw 13730 5.61±$\pm 1.80$ 21.04±$\pm 8.13$ 248 0.62 251 0.58
yo 1226 5.77±$\pm 1.40$ 41.97±$\pm 9.89$ 14 0.48 20 0.64
Table 4: Statistic on the sample used for our analysis (validation and test) for Mozilla Common Voice. Total number of records, average length in seconds and number of tokens as tokenized by Whisper’s pretrained tokenizer, number of male and female speakers, and gini dispersion index of the snippets-per-speaker dispersion, separately by subset.
Lang # records Seconds # Tokens # M Gini (M) # F Gini (F)
cs 2208 9.93±$\pm 5.38$ 60.04±$\pm 31.51$ 45 0.58 16 0.47
de 4064 8.74±$\pm 5.87$ 32.34±$\pm 20.82$ 147 0.55 64 0.57
en 3485 9.92±$\pm 6.17$ 24.48±$\pm 15.05$ 274 0.52 104 0.52
es 3130 11.56±$\pm 6.78$ 42.89±$\pm 25.15$ 95 0.57 48 0.48
fi 1141 9.99±$\pm 5.44$ 46.65±$\pm 26.61$ 25 0.63 16 0.45
fr 3341 10.23±$\pm 6.15$ 38.19±$\pm 23.92$ 151 0.52 84 0.54
hu 2145 10.47±$\pm 6.05$ 62.48±$\pm 35.59$ 45 0.48 21 0.50
it 2419 13.00±$\pm 7.19$ 52.78±$\pm 28.99$ 87 0.47 37 0.53
nl 2351 8.16±$\pm 5.23$ 36.63±$\pm 22.88$ 69 0.54 42 0.54
ro 2701 11.29±$\pm 5.75$ 57.31±$\pm 31.54$ 51 0.52 19 0.57
sk 1266 10.33±$\pm 5.55$ 62.52±$\pm 35.16$ 26 0.57 10 0.47
Table 5: Statistic on the sample used for our analysis (validation and test) for VoxPopuli. Total number of records, average length in seconds and number of tokens as tokenized by Whisper’s pretrained tokenizer, number of male and female speakers, and gini dispersion index of the snippets-per-speaker dispersion, separately by subset.
Lang # records Seconds # Tokens # M Gini (M) # F Gini (F)
ar 836 10.83±$\pm 4.08$ 57.52±$\pm 21.70$ 9 0.31 38 0.51
ca 2914 11.87±$\pm 4.00$ 39.93±$\pm 14.87$ 5 0.15 5 0.05
cs 3772 12.05±$\pm 4.45$ 53.38±$\pm 19.93$ 7 0.36 5 0.21
de 3326 12.98±$\pm 6.32$ 37.19±$\pm 13.30$ 4 0.07 4 0.47
en 2832 9.75±$\pm 3.58$ 24.55±$\pm 8.54$ 13 0.33 15 0.28
es 3082 12.16±$\pm 3.80$ 37.88±$\pm 13.67$ 6 0.38 7 0.35
fi 3482 12.74±$\pm 4.43$ 47.38±$\pm 18.27$ 12 0.34 18 0.47
fr 3194 10.25±$\pm 3.58$ 38.80±$\pm 15.06$ 4 0.06 6 0.15
hu 4374 12.09±$\pm 4.12$ 54.90±$\pm 21.68$ 5 0.03 5 0.05
it 3528 14.53±$\pm 5.34$ 40.59±$\pm 14.30$ 6 0.33 6 0.31
ja 2348 12.99±$\pm 3.77$ 49.50±$\pm 18.25$ 6 0.50 16 0.66
nl 2594 9.55±$\pm 3.73$ 40.59±$\pm 14.63$ 8 0.53 6 0.23
pt 3944 12.50±$\pm 4.24$ 35.32±$\pm 12.88$ 5 0.08 4 0.19
ro 4022 10.22±$\pm 3.63$ 48.98±$\pm 18.03$ 5 0.08 7 0.35
ru 3496 11.39±$\pm 3.97$ 42.57±$\pm 16.14$ 13 0.36 11 0.26
sk 2600 11.65±$\pm 3.81$ 53.15±$\pm 19.93$ 8 0.41 10 0.57
sr 2820 10.76±$\pm 3.40$ 59.46±$\pm 21.59$ 4 0.07 6 0.08
sw 3588 14.08±$\pm 4.52$ 49.60±$\pm 17.25$ 18 0.42 17 0.39
yo 2850 16.32±$\pm 5.66$ 73.91±$\pm 34.05$ 7 0.36 6 0.50
Table 6: Statistic on the sample used for our analysis (validation and test) for Fleurs. Total number of records, average length in seconds and number of tokens as tokenized by Whisper’s pretrained tokenizer, number of male and female speakers, and gini dispersion index of the snippets-per-speaker dispersion, separately by subset. Speakers extracted automatically.
Table 4, 6, and 5 reports all collected statistics on CV, Fleurs, and VoxPopuli, respectively.
A.6 Details on Gender Probing
Training and test set for probing experiments are sampled from the concatenation of each language’s validation and test sets for consistency with error rate experiments. We balanced the samples on gender, stratifying on the speaker distribution.
We used standard logistic regression probes as per scikit-learn’s implementation and standard parameters but a larger allowance of steps for convergence (n=1000). For MDL probes, we used logistic regression as the backbone classifier for comparability and operationalized the probe via the online code configuration Voita and Titov (2020). Following previous work (Orgad et al., 2022), we set the dataset slices to (percentage): [0.2, 0.4, 0.8, 1.6, 3.2, 6.25, 12.5, 25, 50, 100].
We probe a total of 1000 distinct positions that correspond roughly to the first tens seconds of each recording (Whisper encodes audio with 25 milliseconds-long rolling frame and stride of 10 milliseconds) (Radford et al., 2022).
A.7 Energy Statement
Experiments were conducted using a private infrastructure on hardware of type A100 PCIe 80GB (TDP of 250W) with carbon efficiency of 0.233 kgCO2eq/kWh. [^21] https://app.electricitymaps.com Consumption is mostly due to inference to generate transcripts. We estimate a total emission of 3.6 kgCO2eq for experiments in the main body (transcripts of test and validation sets). Transcripts on training sets account for additional 104.24 kgCO2eq. No emission was directly offset. Total emissions are equivalent to driving an internal combustion engine car for 980 km. [^22] Estimate based on average CO2 emissions of new passenger cars in EU in 2020. https://www.acea.auto/figure/average-co2-emissions-of-new-cars-in-eu Estimate was conducted using codecarbon [^23] https://github.com/mlco2/codecarbon but does not account for energy for cooling the infrastructure.
Appendix B Artifacts Statement
The following is a list of artifacts we produced in this work that we release to facilitate future research.
  • Transcriptions obtained with Whisper and SeamlessM4T of the full datasets Mozilla Common Voice, Google Fleurs, and Meta VoxPopuli.
  • Per-record annotations of segments with voice activity extracted from the voice activity detection pipeline (§A.2).
  • Per-record acoustic features in our phonetic study (§5).
  • Per-record embeddings and cluster IDs extracted on Fleurs with the speaker identification pipeline (§A.3).
  • Extensive statistics on speaker, gender, and record distributions on the three datasets—which we could not find anywhere else online.
  • Dataset samples on which we computed the quality and gap metrics for reproducibility purposes (§3).
(a) Probes trained on untouched labels.
(b) Probes trained on shuffled labels.
Figure 9: F-M gender probing F1 Macro performance for every contextual embedding in Whisper-large-v3 within the first 10 seconds (x axis). Logistic regression probe with L2 regularization trained on standard (8(a)) and shuffled (8(b)) training labels. Orange lines indicate the average length of test segments. All CV languages.